OpenStudy (anonymous):

can some one explain to me how to find zeros of the function f(x)= (x+4)^3

OpenStudy (phi):

if you know a*b=0 then you know a=0 or b=0 (or both) here you have (x+4)(x+4)(x+4)=0 when is this zero?

OpenStudy (anonymous):

i don't understand?

OpenStudy (phi):

one step at a time: when is (x+4)= 0 ?

OpenStudy (anonymous):

what do you mean when is?

OpenStudy (phi):

x can be any number. when it is a particular number (x+4)=0 what x makes (x+4) zero?

OpenStudy (anonymous):

Alright Sara When you find the zeroes, you're looking for your x intercepts. Start with this equality f(x)=y (a function with variable x (for example 2x+2) = y what do we know about x intercepts and their y component

OpenStudy (anonymous):